OpenStudy (anonymous):

Which equation represents the solution to the following? I have some lemons. Jill has two more limes than five times the number of lemons that I have. Emma has one-third as many limes as Jill. Emma has 4 limes. How many lemons do I have?

OpenStudy (anonymous):

@aric200

OpenStudy (aric200):

There are 2 questions asked here?

OpenStudy (aric200):

@WhaleShark155

OpenStudy (aric200):

Which equation represents the solution to the following?

OpenStudy (aric200):

How many lemons do I have?

OpenStudy (anonymous):

\[5L+2=\frac{ 4 }{ 3 } , \frac{ 5L+2 }{ 3 }=4 , \frac{ 5L }{ 3} +2= 4, \frac{ 5 }{ L } + \frac{ 1 }{ 3 } + 2 = 4\] are the answer choices

OpenStudy (aric200):

So there is you, Jill, and Emma

OpenStudy (anonymous):

Yes

OpenStudy (aric200):

Ima solve the lemon one first haha

OpenStudy (anonymous):

Okay. I'll try solving it too?

OpenStudy (aric200):

Aye sure sounds good yes.

OpenStudy (anonymous):

So to solve the number of lemons would the equation be (5*2)L/3?

OpenStudy (anonymous):

5(2)L/3=4 I think?

OpenStudy (aric200):

Ima break it down simple as I can.

OpenStudy (anonymous):

Okay and I will listen as best as I can :D

OpenStudy (aric200):

We need to figure out all 3s Totals

OpenStudy (aric200):

We already know Emma has 4 lemmons

OpenStudy (aric200):

And that is 1/3 of Jills total

OpenStudy (aric200):

So 4x3

OpenStudy (anonymous):

Is 12

OpenStudy (aric200):

Jill has 12 Yes

OpenStudy (aric200):

Now the hard part

OpenStudy (anonymous):

Okay :D

OpenStudy (aric200):

Well sorta

OpenStudy (aric200):

We just have to find out how many lemons you have :O Jill has two more limes than five times the number lemons that I/you have.

OpenStudy (anonymous):

So x= 5*(12-2)?

OpenStudy (aric200):

I dont think so I believe It is saying Jill is going to have more lemons

OpenStudy (aric200):

12-2/5

OpenStudy (aric200):

I believe

OpenStudy (aric200):

Im going to verify this

OpenStudy (aric200):

@kropot72 Please assist

OpenStudy (anonymous):

Okay :P

OpenStudy (aric200):

I believe it would leave you with 2 lemons

OpenStudy (anonymous):

How would it be only 2 lemons D:?

OpenStudy (aric200):

But the way its worded kinda got me thinking multiple things

OpenStudy (aric200):

Because she has 2 more then you when your number is multiplied times 5

OpenStudy (aric200):

So you have 2 lemons and * that by 5 you get 10 And she has 2 more

OpenStudy (anonymous):

Yes but I got 11 :I?

OpenStudy (anonymous):

Oh

OpenStudy (aric200):

How did you get 11...

OpenStudy (aric200):

That could not even be possible...

OpenStudy (aric200):

Well in less Jill does not even have 12 lemons

OpenStudy (anonymous):

well I did 12-2/5 and 2 divided by 5 is .4 and 12-.4 is 11.6 and I changed it to 11?

OpenStudy (aric200):

why did you do 2/5?

OpenStudy (anonymous):

Cause you said to

OpenStudy (aric200):

You already did here 12-2/5

OpenStudy (aric200):

You then did it again?

OpenStudy (anonymous):

I guess?

OpenStudy (aric200):

Kropot72 Save me haha

OpenStudy (anonymous):

I thought a human calculator knows best? :o

OpenStudy (aric200):

I feel as if i know but you got me second guessing <.<

OpenStudy (anonymous):

Oops :3

OpenStudy (aric200):

Im a human calculator when it comes to 2 + 2

OpenStudy (anonymous):

hahaha

OpenStudy (kropot72):

Let x be the number of lemons that I have. The number of limes that Jill has = 5x + 2 The number of limes that Emma has is give by: \[\frac{5x+2}{3}=4\ .........(1)\] Multiplying both sides of (1) by 3 we gwt: 5x + 2 = 12 5x = 10 x = 2

OpenStudy (kropot72):

given*

OpenStudy (anonymous):

Goodness your smart o.o

OpenStudy (aric200):

Okay so we both got 2 :D

OpenStudy (kropot72):

Cool!

OpenStudy (aric200):

Just wanted to verify

OpenStudy (aric200):

I feel as if that is a more complex way of doing it though.

OpenStudy (kropot72):

Sure. Usually the fastest method is best, although the fastest way might not be the easiest to follow.

OpenStudy (anonymous):

Why is math so difficult ;-;

OpenStudy (aric200):

Geometry is way worse so dont even start <.<

OpenStudy (anonymous):

MAKE THE LLAMAS DO THE WORK :D

OpenStudy (kropot72):

As one of my teachers put it: "Math has a naked beauty all its own, once its mysteries are unveiled."

OpenStudy (aric200):

Your teacher worries me haha

OpenStudy (anonymous):

I have to go soon :(

OpenStudy (anonymous):

mady here to help

OpenStudy (aric200):

Mady get out of here your like an hour late :O

OpenStudy (anonymous):

OOH SHOOED BY THE ARIC :P

OpenStudy (anonymous):

well excuse mee

OpenStudy (aric200):

I kid your here now so you can give me a medal <3

OpenStudy (anonymous):

So you gave up on helping me o.o?

OpenStudy (anonymous):

mhm thats what i thought xD

OpenStudy (aric200):

No But if you want me to answer your other one you must repost

OpenStudy (aric200):

Oh thank you for da medals peeps.

OpenStudy (aric200):

2 Chainzzz

OpenStudy (anonymous):

anytime no liee 2 chainzz

OpenStudy (aric200):

2 Medalsss

OpenStudy (anonymous):

Wow your all weirdoz XDD But awesome ^-^

OpenStudy (aric200):

Says you ....

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

thanks! <3 ik im awesome (;

OpenStudy (anonymous):

ha

OpenStudy (aric200):

Everyone is weird in there own ways :D

OpenStudy (anonymous):

BUT LLAMAS ARE REALLY COOL AND FLUFFY. :P

OpenStudy (aric200):

Yeah I thought about dating one.

OpenStudy (anonymous):

they spit on people dont they .. ewy

OpenStudy (aric200):

but there cool and im just like :( a loser ya know

OpenStudy (anonymous):

I bet that would be a beautiful couple o.o

OpenStudy (anonymous):

pshh pleasee aric

OpenStudy (aric200):

Im in online school I mean I am such a swegster ya know

OpenStudy (aric200):

gettin all the ladies

OpenStudy (aric200):

So many friends

OpenStudy (anonymous):

I have to leave soon :(